could you please help me find the answer to this ? Find the slope , if it exists of the line containing the pair of points (5,9) and (8,-2)
Answer by mananth(16946)

The slope of line passing through (x1,y1),(x2,y2) = (y2-y1)/(x2-x1)
(5,9) and (8,-2)
m= -2-9 / 8-5
m=-11 / 3
slope = -11/3